INAF queue observing, 2019 April 30-May 1

Observer: F. Cusano, A. Gargiulo
Telescope Operator: D. Gonzalez-Huerta
Support Astronomer: O. Kuhn
Instrument:

At the sunset the sky is clear.

Summary

Night duration (12 degree el twilight) 3:00 - 11:35 UT = 8 h 35m

Observing time: 6*h* 43*min* 78(%)

Weather time loss: 0*h* 0*min* 0(%)

technical time loss 1*h 52*min* 22(%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibrations flat and arc for Gilli mask 47 MODS1/MODS2 Dual     mods1b.0009.fits not good
Gilli mask 47 MODS1/MODS2 Dual 4.66/4h mods1b(r).11-17(15-21), mods2b(r).11-16(15-21) completed mods1b.11, mods1r.15, mods2b.11, mods2r.15 not good
Marziani HIP71172 LUCI1/2 G200-HK     do not use luci2.0005
Marziani SDSSJ14 LUCI1/2 G200-HK 0.7/0.7h luci1.0048-0053 luci2.0013-0018 completed
Bianchi HIP87643 LUCI2/ESM G210-J      
Bianchi MRK695 LUCI2/ESM G210-J 0.9/2h luci2.0032-0042 completed
Rossi 070521 LUCI2/ESM G200-HK 0.7/4h luci2.0047-0061  
Rossi HIP87643 LUCI2 G200-HK      

Details

Calibrations
  • flat for Gilli mask 47 mods1b.20190501.0002-0007 mods1r.20190501.0002-0007 mods2b.20190501.0003-0007 mods2r.20190501.0002-0007
  • arc for Gilli mask 47 mods1b.0008-0010 mods1r.0008-0010 mods2b.0008-0010 mods2r.0008-0010 file mods1b.20190501.0009.fits not good

02:55 seeing on dimm 2.27"

02:59 trying to slew to Gilli mask 47: this is the last mask to close the program. Tomorrow, one of the two MODS will be dismissed, and the day after tomorrow both MODS will be out. In the next run the field will be not visible anymore. The seeing is higher than required, but the trend seems to go toward better conditions, so we decide to try since this is our last chance to close the program in this year.
  • 03:04 seeing on dimm 1.7"
  • 03:04 acq mods1r.0011-0014 mods2r.0011-0014
  • 03:16 seeing on dimm jump to 3.0"
  • 03:24 seeing on dimm 1.2"
  • 03:24 sending science
  • 03:33 seeing on dimm jump to 2"
  • 03:40 seeing on dimm 1.7"
  • 03:55 seeing on dimm 1.37" , on guider 1.14"
  • 03:58 seeing on dimm 1.2"
  • 04:00 seeing on dimm 1.6"
  • 04:08 seeing on dimm 1.2"
  • 04:11 seeing on dimm 0.99"
  • 04:15 seeing on dimm 0.87"
  • 04:26 seeing on dimm 0.8"
  • 04:32 seeing on dimm 0.6"
  • 04:51 seeing on dimm 0.67"
  • 05:15 seeing on dimm 0.69"
  • 05:23 seeing on dimm 1.2"
  • 05:33 seeing on dimm 1.2"
  • 05:39 seeing on dimm 1.3"
  • 05:43 seeing on dimm 1.3"
  • 05:48 seeing on dimm 1.0"
  • sci mods1b.0011-0017 mods1r.0015-0021 mods2b.0011-0016 mods2r.0015-0021 mods1b.11, mods1r.15, mods2b.11, mods2r.15 no good, seeing too variable; mods2b has one less image for telescope problem

06:00 we switch to LUCI

06:42 we slew to Marziani's Telluric HIP71172
  • acq luci1.0037-0039 luci2.0001-0003
  • sci luci1.0040-0043 ==luci2.0005--0008 do not use luci2.0005
  • seeing on dimm 1.1"

06:58 we slew to Marziani's target SDSSJ14
  • acq luci1.0044-0047 luci2.0009-0012
  • 07:15 seeing on dimm 0.9"
  • sci luci1.0048-0053 luci2.0013-0018

07:39 wes slew to Bianchi's Telluric HIP87643
  • 07:42 mos error
  • 07:56 mos error solved
  • 07:57 seeing on dimm 1.2"
  • 07:59 mos error
  • 08:04 mos error solved
  • 08:05 mos error again
  • 08:10 seeing on dimm 1.03"
  • 08:12 problem solved
  • acq luci2.0019-0021
  • 08:16 seeing on dimm 1.1"
  • sci luci2.0022-0025

08:26 we slew to Bianchi's target MRK695
  • 08:35 collimation problem
  • 08:36 adsec failure
  • 08:56 we start integration
  • acq luci2.0026-0031
  • 09:00 seeing on dimm 1.2"
  • 09:35 seeing on dimm 0.87"
  • sci luci2.0032-0042

10:00 slew to Rossi's target 070521
  • 10:03 mos error
  • 10:16 control panel stucks, we re-open it
  • 10:19 we send the script again
` * 10:20 softwere problems
  • 10:33 we start acquisition
  • 10:41 seeing on dimm 0.72"
  • acq luci2.0043-0046
  • 10:51 seeing on dimm 0.85
  • 11:33 seeing on dimm 0.8"
  • sci luci2.0047-0061

11:35 we slew to Rossi's Telluric HIP87643
  • 11:37 check pointing
  • acq luci2.0062-0064
  • sci luci2.0065-0068

Calibrations
  • Flat for Marziani luci1.0054-0058 luci2.0069-0073
  • Arc for Marziani luci1.0059-0069 luci2.0074-0084
  • Flat for Gilli mask 47 mods1b.0018-0023 mods1r.0022-0027 mods2b-0017-0022 mods2r.0022-0027
  • Arc for Gilli mask 47 mods1b.0024-0026 mods1r.0028-0030.fits mods2b.0023-0025 mods2r.0028-0030
Topic revision: r7 - 02 May 2019, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback